# Optical and UV Flares from Binary Black Hole Mergers in Active Galactic Nuclei

### v1

@@ III.2 Optical light-curves
Figure 10: Examples of the flare time lag and its components as a function of the merger location $a$. These periods are calculated in the observer frame. The curves initiate when condition 37 is fulfilled. The parameter configurations for the curves in each panel are specified in the plot labels and in the text.
IV Optical counterpart candidates to BBH mergers
Graham et al. [11, 13] have proposed a number of optical flares measured by the ZTF from known AGNs as possible EM counterparts associated with LVK BBH merger events. These counterpart candidates are reported with time delays $\Delta t_{\ell}\lesssim 200$ days after the associated GW events, in AGNs situated at $z>0.3$ with central engines of $M_{\mathrm{S}}>10^{8}$ M. Such MM associations are currently under debate [15, 16, 17], mostly because of the large spatial error of the LVK events. Most of these candidate flares are not compatible with the present MM scenario, which predict optical flares with $\Delta t_{\ell}\lesssim 200$ days provided $M_{\mathrm{S}}<10^{7}$ M, and such flares could be detected up to $z\sim 3$ with the sensitivity of the ZTF. In AGNs with $M_{\mathrm{S}}\gtrsim 10^{8}$ M, the flare delays would be likely of year scales under the MM scenario discussed here (see Figure 8). Thus, if the aforementioned candidate flares are real EM counterparts, they would be better explained by mechanisms with faster interaction among the BBH remnant and the AGN disc, such as the re-directed jet scenario discussed in [23, 50].
On the other hand, more precise measurements for the properties of the aforementioned AGN candidates could change the scenario that better explain the potential MM association. For concreteness, we note the mass of the central engine in the AGN J053408.41+085450.6 is currently not accurately measured as pointed out in [13]. Furthermore, J053408.41+085450.6 is reported with photometric redshift, which is a less accurate method compared to spectroscopic redshift measurements of the other AGN candidates. If upgraded measurements lead to an SMBH mass close to $\sim 10^{7}$ M and a redshift $z\lesssim 0.3$ for J053408.41+085450.6, the optical flare peaking at $\sim$ MJD 58900 from this source could be naturally explained by the present emission model as the counterpart to the GW event with ID GW190731_140936. We leave such investigation as the subject of a forthcoming work.
V Summary and conclusions
@@
In this paper, we present a physical, analytic model that predicts the time delay, duration, spectrum, and optical LC of EM feedback from stellar mass BBH mergers within thin AGN discs. Such environments have been proposed as promising hosts for the observed GW events associated to BBH mergers measured by the LIGO/Virgo/Kagra experiment. The EM processes considered come from the thermal emission of plasma eruptions expanding outside the disc, expelled from the disc by the BH merger remnant. In our analytic approach (see section II for details), we assume that the BBH is embedded in the disc and co-rotating. After the merger, the remnant is gravitationally kicked out, accretes from the unperturbed disc and produces a quasi-spherical outflow which expels disc material. These ejections expand and cool outside the AGN disc, producing thermal flares which we model here.
@@
In our formulation, the EM counterpart is determined by nine parameters characterising the hosting AGN ($D_{\mathrm{L}}$, $M_{\mathrm{S}}$, $\dot{M}_{\mathrm{S}}$), the BBH merger ($a$, $M_{\bullet}$, $v_{\mathrm{k}}$, $\theta_{\mathrm{k}}$), and the remnant’s outflow ($\eta_{\mathrm{in}}$, $\eta_{\mathrm{w}}$; see Section II for details). When the model is applied to interpret a particular LC dataset, MM observations can constrain the parameters $D_{\mathrm{L}}$, $M_{\mathrm{S}}$, $\dot{M}_{\mathrm{S}}$, and $M_{\bullet}$, and then the possible LCs predicted by the model would depend on the other five parameters.
@@
    The flare emission can be comparable to or outshine that of the hosting AGN at UV-optical frequencies, for mergers occurring in AGNs with SMBHs of $10^{6-7}$ M.
@@
    The kick angle relative to the disc plane $\theta_{\mathrm{k}}$ is a crucial parameter that modulates the flare profile and determines whether EM feedback is possible (see e.g., equations 10 and 37). Within the parameter space discussed here, our results suggest that significant EM counterparts are possible when $\theta_{\mathrm{k}}\lesssim 10^{\circ}$.
@@
    In AGNs with SMBHs of $\sim 10^{6}(10^{7})$ M, distinguishable flares have time delays of a few tens (hundreds) of days and require remnants with kick velocities $\gtrsim 300$ (100) km s-1. For hosts with SMBHs of $\sim 10^{8}$ M (and heavier), the flares have time lags of thousands of days (and longer). We then suggest that AGNs with SMBHs masses $\lesssim 10^{7}$ M are the most suitable hosts for localising EM counterparts under the present scenario.
@@
    EM counterparts from AGNs with SMBHs of $10^{6-7}$ M are suitable targets for time-domain UV space telescopes like the forthcoming ULTRASAT [85] and UVEX [86], as well as for optical time domain surveys (see Figures 4-5).
@@
    We estimate that ZTF can detect the flares predicted here up to $z\sim 0.04-0.09$ from AGNs with $M_{\mathrm{S}}=10^{6}$ M, and up to $z\sim 0.1-0.3$ when $M_{\mathrm{S}}\sim 10^{7}$ M. Rubin could observe the flares up to $z\sim 0.1-0.3$ and $z\sim 0.5-1$ when $M_{\mathrm{S}}\sim 10^{6}$ and $10^{7}$ M, respectively.
@@ Acknowledgements.
JCRR acknowledges support from Rio de Janeiro State Funding Agency FAPERJ, grant E-26/205.635/2022. CdB acknowledges the financial support from CNPq (316072/2021-4) and from FAPERJ (grants 201.456/2022 and 210.330/2022) and the FINEP contract 01.22.0505.00 (ref. 1891/22). RN acknowledges support from CNPq through a Bolsa de Produtividade and NASA through the Fermi Guest Investigator Program (Cycle 16). The authors made use of Sci-Mind servers machines developed by the CBPF AI LAB team and would like to thank P. Russano and M. Portes de Albuquerque for all the support in infrastructure matters.
Appendix A The initial pressure of the plasma ejection

### v2

@@ III.2 Optical light-curves
Figure 10: Examples of the flare time lag and its components as functions of the merger location $a$. These periods are calculated in the observer frame. The curves begin and end when the condition for the flux ratio $r_{\nu}>0.25$ (see equation 38) and the condition 37 are no longer satisfied. The parameter configurations for each panel are given in the plot labels and the text.
IV Flare delay, kick velocity, and the $q-\chi_{\mathrm{eff}}$ correlation
The present MM emission model can be used to investigate constraints on MM observables, such as the binary effective spin $\chi_{\mathrm{eff}}$, the mass ratio $q=m_{2}/m_{1}$, the merger remnant mass $M_{\bullet}$, and the time delay $\Delta t_{\mathrm{\ell}}$ of the EM signature. In this formulation, the flare delay $\Delta t_{\ell}$ can be linked to the BBH effective spin by parametrising the remnant’s kick velocity $v_{\mathrm{k}}$ and direction $\theta_{\mathrm{k}}$ (relative to the disc plane; see Figure 1) in terms of the dimensionless spins $\bar{\chi}_{1}$, $\bar{\chi}_{2}$, and the masses $m_{1}$, $m_{2}$ of the merging BHs. To achieve this, we use the empirical formulae of [58, 59] (see also [93]), which provide analytic fits to numerical relativity simulations of BBH merger recoils.
In terms of the kick velocity components, the kick angle relative to the disc plane is given by:
$\theta_{\mathrm{k}}=\arcsin\left\{\frac{v_{||}}{|\vec{v}_{\mathrm{k}}|}\right\},$ (42)
where $|\vec{v}_{\mathrm{k}}|$ is the kick velocity magnitude, and $v_{||}$ is its component normal to the disc plane. Assuming the BBH orbital plane is aligned with the AGN disc plane, the Cartesian components of the recoil velocity $(v_{x},v_{y},v_{\parallel})$ can be expressed in terms of the spin vectors $\vec{\chi}_{1},\vec{\chi}_{2}$ and the mass ratio $q=m_{2}/m_{1}\leq 1$ as given by [58, 59]:
$v_{\mathrm{||}}=\frac{16\eta^{2}}{1+q}\left[v_{0}+v_{A}S_{||}+v_{B}S_{||}^{2}+v_{C}S_{||}^{3}\right]$ (43)
×$\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\times|\vec{\chi}_{\perp,2}-q\vec{\chi}_{\perp,1}|\cos\left(\delta_{\triangle}-\delta_{1}\right),$
$v_{\mathrm{x}}=v_{m}+\cos\xi v_{\perp},$ (44)
$v_{\mathrm{y}}=v_{\perp}\sin\xi,$ (45)
with
$v_{m}=A\eta^{2}\sqrt{1-4\eta}(1+B\eta),$ (46)
$v_{\perp}=\frac{H\eta^{2}}{1+q}\left(\chi_{||,2}-q\chi_{||,1}\right),$ (47)
$\eta=\frac{q}{(1+q)^{2}},$ (48)
where we adopt the following coefficient values: $A=1.2\times 10^{4}$ km/s, $B=-0.93$, $H=6.9\times 10^{3}$ km/s, $v_{0}=3678$ km/s, $v_{A}=2481$ km/s, $v_{B}=1793$ km/s, $v_{C}=1507$ km/s, and $\xi=145^{\circ}$, as given in [59] and references therein. In the equations above, $\vec{\chi}_{\perp,1}$ and $\vec{\chi}_{\perp,2}$ are the projections of the spin vectors onto the orbital plane (i.e., in-plane components), and $S_{\parallel}$ is the component parallel to the binary’s orbital angular momentum of the vector
$\vec{S}=2\frac{\vec{\chi}_{2}+q^{2}\vec{\chi}_{1}}{(1+q)^{2}}.$ (49)
In equation 43, $\delta_{\triangle}$ is the angle between the in-plane component of
$\vec{\Delta}=\frac{(m_{1}+m_{2})^{2}}{(1+q)}(\vec{\chi}_{2}-q\vec{\chi}_{1}),$ (50)
and the in-fall direction at merger, while $\delta_{1}$ is an angular phase dependent on the binary’s initial separation.
We define the $x$-direction of the spin vectors and kick velocity as parallel to the in-fall direction at merger and set $\delta_{1}=0$. The choice of in-plane coordinates does not affect the emission outcome discussed in Section II, as we assume the local environment around the merger to be azimuthally symmetrical. Finally, we parameterise the spin vector components as:
$\chi_{\parallel,i}=\chi_{i}\cos\varphi_{i},$ (51)
$\chi_{x,i}=\chi_{i}\sin\varphi_{i}\cos\phi_{i},$ (52)
$\chi_{y,i}=\chi_{i}\sin\varphi_{i}\sin\phi_{i},$ (53)
where $\varphi_{i}$ and $\phi_{i}$ are the polar and azimuthal angles, $\chi_{i}\equiv|\vec{\chi}_{i}|$ are the dimensionless spin magnitudes, and $i=1,2$ label the heavier and lighter BHs, respectively. Thus, the mass ratio $q$ and effective spin $\chi_{\mathrm{eff}}$ of a BBH with spins $\vec{\chi}_{i}$ and masses $m_{i}$ are given by:
$q=\frac{m_{2}}{m_{1}}\leq 1,\,\,\,\,\,\chi_{\mathrm{eff}}=\frac{\chi_{1}\cos\varphi_{1}+q\chi_{2}\cos\varphi_{2}}{1+q}\in[-1,1].$ (54)
Summarising, the observables ($\chi_{\mathrm{eff}},q,M_{\bullet},v_{\mathrm{k}},\theta_{\mathrm{k}},\Delta t_{\ell}$) can be predicted using the MM emission model from Section II along with equations (42)-(54), for a given configuration $(m_{i},\varphi_{i},\phi_{i},\chi_{i})$ of the BBH components.
BBH mergers detected by the LVK experiment show a trend where the effective spin $\chi_{\mathrm{eff}}$ increases towards positive values as the $q$ ratio decreases [41, 94, 95]. This suggests a bias in the AGN GW channel towards alignment of the heavier BH’s spin with the orbital angular momentum (i.e., $\varphi_{1}\sim 0$, see equation 54), possibly due to torques exerted by gas accretion.
In Figures 11-12, we show the distributions of kick velocities and directions, and their associated configurations in the $q-\chi_{\mathrm{eff}}$ space, resulting from imposing the $\varphi_{1}=0$ constraint on the recoil velocity model of equations (43)-(45). We consider three scenarios for the spin of the lighter BH: (i) a uniform angular distribution, (ii) a quasi-aligned spin, and (iii) a quasi-anti-aligned spin relative to the orbital angular momentum. To emulate these cases, we keep $\varphi_{1}=0$ and sample $10^{4}$ parameter configurations within the intervals $q\in[0.05,1]$, $\chi_{1,2}\in[0.05,1]$, $\phi_{1,2}\in[0,2\pi]$, with the spin of the lighter BH sampled as:
  1. (i)
    $\varphi_{2}\in[0,2\pi]$ (uniform distribution),
  2. (ii)
    $\varphi_{2}\in[0,\delta\varphi]$ (quasi-aligned),
  3. (iii)
    $\varphi_{2}\in[\pi-\delta\varphi,\pi]$ (quasi-anti-aligned).
where $\delta\varphi=\pi/10$.
The $v_{\mathrm{k}}-\theta_{\mathrm{k}}$ and $q-\chi_{\mathrm{eff}}$ correlations resulting from the three scenarios above are presented in the plots in different columns of Figure 11, as indicated. Among these, we identify that binaries with quasi-anti-aligned spins (right column in Figure 11) are the most likely to produce post-merger EM counterparts under the emission scenario discussed in this paper. This spin configuration concentrates most of the kick orientations quasi-parallel to the AGN disc plane (as shown in Figure 12), while also being consistent with the observed $q-\chi_{\mathrm{eff}}$ correlation. The $\varphi_{1}\in[0,\pi]$ case produces relatively few kicked remnants quasi-parallel to the disc, whereas the quasi-aligned case does not produce null $\chi_{\mathrm{eff}}$ for $q\rightarrow 1$, as observed.
The $q-\chi_{\mathrm{eff}}$ relation provides an additional constraint that can be used to favour or rule out candidate EM counterparts under the MM emission scenario discussed in this paper. We illustrate this by deriving the observables ($\chi_{\mathrm{eff}}$, $q$, $M_{\bullet}$, $v_{\mathrm{k}}$, $\Delta t_{\ell}$) for EM signatures in the optical $g$-band, restricted to flux ratios $r_{\nu}>0.25$ (see Section III), produced by BBH mergers within a SG03 disc, where the BBH merger follows configuration (iii) above (the case of BHs with anti-aligned spins). Given the values for the SMBH mass $M_{\mathrm{S}}$ and its normalised accretion rate $\dot{m}_{\mathrm{S}}$, we sample $10^{4}$ parameter configurations of BBH mergers within the following ranges:
  • $a\in[300R_{\mathrm{g}},a_{\mathrm{max}}]$,
  • $m_{1}\in[5,80]$,
  • $q\in[0.05,1]$,
  • $\chi_{1},\chi_{2}\in[0.05,1]$,
  • $\phi_{1},\phi_{2}\in[0,2\pi]$,
  • $\varphi_{2}\in[\pi-\delta\varphi,\pi]$.
with $\varphi_{1}$ fixed at 0. The first range listed corresponds to the radial location of the merger within the AGN disc, with the upper limit determined by the outer boundary of the disc (as specified by the criterion in equation 55). Since $a$ spans several orders of magnitude, we sample it uniformly on a logarithmic scale. As in the example leading to Figures 8 and 9, here we consider the AGN situated at $D_{\mathrm{L}}=1600$ Mpc and fix the efficiencies related to the remnant’s wind to $\eta_{\mathrm{in}}=0.1$ and $\eta_{\mathrm{w}}=0.05$.
The predictios for ($\chi_{\mathrm{eff}}$, $q$, $M_{\bullet}$, $v_{\mathrm{k}}$, $\Delta t_{\ell}$) are displayed in Figure 13. From these outcomes, we observe that, for instance, visible EM counterparts with time lags of $\Delta t_{\ell}\sim 10$ days can only be produced in hosts with SMBHs of $\sim 10^{7}$ Mdirect-$\odot$ or smaller. Such EM signatures are unlikely to be produced by mergers with $q\lesssim 0.2$, $\chi_{\mathrm{eff}}\gtrsim 0.6$, or remnants of $M_{\bullet}\gtrsim 40$ Mdirect-$\odot$. Mergers with these values of $q$, $\chi_{\mathrm{eff}}$, and $M_{\bullet}$ in hosts with $M_{\mathrm{S}}\lesssim 10^{7}$ M would produce flares with time delays of $\Delta t_{\ell}\gtrsim 50$ days. In low-mass AGNs, the EM signature is more likely to be produced when the AGN has a relatively high accretion rate, whereas low accretion rates favour EM emission in high-mass AGNs, according to Figure 13. We also observe that, in general, the kick velocity must be $\gtrsim 150$ km s-1 to produce flares with time delays of $\Delta t_{\ell}<50$ days.
Figure 11: Upper: Merging BBHs in the $q-\chi_{\mathrm{eff}}$ space, that results from sampling $10^{4}$ configurations of spin and masses of the binary components (see the text). Lower: Correlation of the kick magnitude and its direction relative to the AGN disc plane, associated to the sampled configurations of the upper panels. The kick velocity and direction are obtained adopting an analytical parametrisation presented in [58, 59], and detailed in the text. All panels were obtained assuming the spin of the heavier BH aligned with the orbital spin, whereas in the left, centre, and right panels, the polar angle of the lightest BH spin, $\varphi_{2}$ is constrained as indicated.
Figure 12: Distribution of the kick velocity magnitudes (upper) and directions relative to the disc plane (lower) associated to the lower panels of Figure 11.
Figure 13: Correlation of the optical flare delay $\Delta t_{\ell}$ with the properties of the BBH merger as plotted in the different columns. The associations represented with the markers are derived from the $10^{4}$ parameter configurations sampled as described in the text. The plots in different rows are obtained considering different masses for the SMBH host, as indicated. Different colours and orientation of the markers corresponds to different AGN accretion rates and flare directions, (see section II), as indicated.
V Optical counterpart candidates to BBH mergers
Graham et al. [11, 14] have proposed a number of optical flares measured by the ZTF from known AGNs as possible EM counterparts associated with LVK BBH merger events. Such MM associations are currently under debate [15, 17, 18], mostly because of the large spatial error of the LVK events.
Among the proposed counterparts, the flare beginning at MJD $\sim$58650 from the AGN J124942.3+344929, located at $z=0.438$, with a central SMBH mass estimated within $10^{8-9}$ M [11] is perhaps the most extensively discussed candidate. This flare has been associated with the event GW190521, with inferred $\chi_{\mathrm{eff}}=0.08$ and $q=0.8$, producing a remnant BH of $\sim 150$ M placing it in the category of an intermediate-mass black hole. The claimed EM counterpart increased the source’s emission by 0.3 mag, peaking approximately 50 days after GW190521.
This flare marginally falls within the parameter space of visible flares derived in Figure 13. As shown in the figure, remnants of $\sim 150$ Mdirect-$\odot$ in AGNs with $M_{\mathrm{S}}\sim 10^{8}$ Mdirect-$\odot$ produce flares with time lags of $\Delta t_{\ell}\gtrsim 50$ days. Furthermore, since J124942.3+344929 is more distant than $D_{\mathrm{L}}=1600$ Mpc (the distance assumed in Figure 13), and its SMBH may be more massive than $10^{8}$ M [11], the predicted time delays from our model could be even longer unless more optimistic parameters are considered.
On the other hand, the EM counterparts proposed in [14] originate from AGNs at $z>0.3$ with central SMBHs of $M_{\mathrm{S}}>10^{8}$ M and exhibit time lags of up to $\sim 200$ days after their potential GW associations. These flares align more comfortably with the parameter space of visible flares presented in Figure 13.
A detailed investigation of these EM candidates requires an appropriate model for the baseline emission (i.e., the source’s emission before and after the flare) for each individual AGN. Such considerations can clearly alter the predictions obtained here, where a generic SED template was used to represent the host’s background emission (see equation 39). Therefore, we leave a comprehensive analysis of the flares proposed in [11, 14] under the MM emission scenario presented here for future work.
VI Summary and conclusions
@@
In this paper, we present an analytic model that predicts the time delay, duration, spectrum, and optical LC of EM feedback from stellar mass BBH mergers within thin AGN discs and the associated mass ratio and effective spin of the merger. AGNs have been proposed as promising hosts for the observed GW events associated to BBH mergers measured by the LIGO/Virgo/Kagra experiment. The EM processes considered come from the thermal emission of plasma eruptions expanding outside the disc, expelled from the disc by the BH merger remnant. In our analytic approach (see section II for details), we assume that the BBH is embedded in the disc and co-rotating. After the merger, the remnant is gravitationally kicked out, accretes from the unperturbed disc and produces a transient quasi-spherical outflow which expels disc material. These ejections expand and cool outside the AGN disc, producing thermal flares which we model here.
@@
In our formulation, the EM counterpart is determined by nine parameters characterising the hosting AGN ($D_{\mathrm{L}}$, $M_{\mathrm{S}}$, $\dot{M}_{\mathrm{S}}$), the BBH merger ($a$, $M_{\bullet}$, $v_{\mathrm{k}}$, $\theta_{\mathrm{k}}$), and the remnant’s outflow ($\eta_{\mathrm{in}}$, $\eta_{\mathrm{w}}$; see Section II for details). When the model is applied to interpret a particular LC dataset, MM observations can constrain the parameters $D_{\mathrm{L}}$, $M_{\mathrm{S}}$, $\dot{M}_{\mathrm{S}}$, $M_{\bullet}$, and $v_{\mathrm{k}}$ (through $q$ and $\chi_{\mathrm{eff}}$), and then the possible LCs predicted by the model would depend on the other four parameters.
@@
    The flare emission can be comparable to or outshine that of the hosting AGN at UV-optical frequencies, for mergers occurring in AGNs with SMBHs of $10^{6-8}$ M.
@@
    The kick angle relative to the disc plane $\theta_{\mathrm{k}}$ is a crucial parameter that modulates the flare profile and determines whether EM feedback is possible (see e.g., equations 13 and 37). Within the parameter space discussed here, our results suggest that significant EM counterparts are possible when $\theta_{\mathrm{k}}\lesssim 10^{\circ}$. Our results suggest that binaries with quasi-anti-aligned spins are the most likely to produce EM counterparts. This spin configuration concentrates most of the kick orientations at small angles relative to the AGN disc plane while also aligning with the observed $q-\chi_{\mathrm{eff}}$ correlation trend.
@@
    In AGNs with SMBHs of masses $\lesssim 10^{7}$ Mdirect-$\odot$, mergers can produce distinguishable flares with time delays as short as $\sim$10 days, requiring kick velocities of $\gtrsim 200$ km s-1. For hosts with SMBHs of $\sim 10^{8}$ Mdirect-$\odot$, flares with time lags of a few tens of days are possible.
@@
    EM counterparts from AGNs with SMBHs of $10^{6-8}$ M are suitable targets for time-domain UV space telescopes like the forthcoming ULTRASAT [96] and UVEX [97], as well as for optical time domain surveys (see Figures 4-5).
@@
    We estimate that ZTF can detect the flares predicted here up to $z\sim 0.04-0.1$ from AGNs with $M_{\mathrm{S}}=10^{6}$ M, and up to $z\sim 0.1-0.3$ when $M_{\mathrm{S}}\sim 10^{7}$ M. Rubin could observe the flares up to $z\sim 0.25-0.5$ and $z\sim 0.7-1.2$ when $M_{\mathrm{S}}\sim 10^{6}$ and $10^{7}$ M, respectively.
@@ Acknowledgements.
We are thankful to the anonymous reviewer for providing valuable suggestions that improved the quality of this work. JCRR acknowledges support from Rio de Janeiro State Funding Agency FAPERJ, grant E-26/205.635/2022. CdB acknowledges the financial support from CNPq (316072/2021-4) and from FAPERJ (grants 201.456/2022 and 210.330/2022) and the FINEP contract 01.22.0505.00 (ref. 1891/22). RN acknowledges support from CNPq through a Bolsa de Produtividade and NASA through the Fermi Guest Investigator Program (Cycle 16). The authors made use of Sci-Mind servers machines developed by the CBPF AI LAB team and would like to thank P. Russano and M. Portes de Albuquerque for all the support in infrastructure matters.
Appendix A AGN thin disc models
Here, we compare the radial profiles of AGN thin discs as derived from the models of Shakura & Sunyaev ([61]; SS73), Sirko & Goodman ([68]; SG03), and Thompson et al. ([69]; TQM05), focusing on the physical properties relevant to the multi-messenger emission scenario discussed in this paper.
To calculate the SS73 radial profiles, we employ the prescription presented in [98] for the inner, middle, and outer radial regions (according to the gas/radiation pressure and opacity dominance). To obtain the radial profiles of the SG03 and TQM05 models, we employ the pAGN, Python based code [70] [^3] https://github.com/DariaGangardt/pAGN with some additional specifications motivated as follows.
The SG03 disc model predicts an additional SED bump feature at near-infrared compared to the disc spectrum of the standard SS73 disc. This additional emission originates at the disc outer regions, due to the implemented heating mechanism generating radiation pressure against gravitational collapse. This emission excess can be employed to limit the size of the disc, so that the disc emission matches observations. To generate emission from SG03 discs consistent with the host galaxy emission template employed in the present study (see equation 39 of the main text), we set the outer boundary of the SG03 disc as
$a_{\mathrm{max}}=\left(\frac{M_{\mathrm{SMBH}}}{10^{8}\mathrm{M}_{\odot}}\right)^{-1/2}10^{4}R_{\mathrm{g}}.$ (55)
with $R_{\mathrm{g}}=GM_{\mathrm{S}}/c^{2}$.
The TQM05 model addresses gas support against self-gravity in the outer region of the disc through star formation by-products. This formulation requires specifying the relation of the velocity dispersion $\sigma$ with the central mass provided by observations, for which we employ the one obtained by [100] (their equation 7). A distinctive characteristic of the TQM05 disc solution is that the accretion rate changes as a function of radial distance to the central engine due to mass loss from star formation. To produce TQM05 solutions that can be compared to the SS73 and SG03 counterparts, we modulate the accretion rate $\dot{M}\mathrm{out}$ of the TQM05 disc at a distance $r\mathrm{out}=10^{7}R_{\mathrm{S}}$, where $R_{\mathrm{S}}=2GM_{\mathrm{SMBH}}/c^{2}$, to obtain disc solutions with the desired accretion rate $\dot{M}\mathrm{S}$ at $3R\mathrm{S}$ (the innermost stable circular orbit of a Schwarzschild SMBH). Specifically, to produce TQM05 disc solutions with accretion rates of $\dot{M}\mathrm{S}=0.0075$, $0.0100$, $0.0250$, $0.0500$, and $0.1000$ $\dot{M}\mathrm{Edd}$, we set the outer accretion rates to $\dot{M}\mathrm{out}=0.450$, $0.920$, $1.972$, $2.925$, and $3.700$ $\dot{M}\mathrm{Edd}$ for an SMBH of $10^{8}$ Mdirect-$\odot$, and $\dot{M}\mathrm{out}=0.0075$, $0.0100$, $0.0290$, $0.1100$, $0.2970$ $\dot{M}\mathrm{Edd}$ for an SMBH of $10^{7}$ Mdirect-$\odot$. For SMBHs of $10^{6}$ Mdirect-$\odot$, our calculations require $\dot{M}\mathrm{out}=\dot{M}_{\mathrm{S}}$.
In Figure 14, we illustrate selected disc properties obtained from the SS73, SG03, and TQM05 models as specified above, and in Figure 15, we show their predicted spectra for different configurations of SMBH masses and accretion rates, as indicated. We observe that, within the radial domain shown in Figure 14, the SG03 model predicts properties of the same order of magnitude as those of the SS73 model. The curves produced by the SG03 model terminate at the outer disc boundary given by equation 55. The TQM05 model, on the other hand, predicts gas densities that are 2–3 orders of magnitude lower than those of the SG03 and SS73 models and a disc aspect ratio about one order of magnitude lower for SMBHs of $10^{7-8}$ M. The spectra emitted by these disc models converge in the optical and UV bands to the template emission for the host galaxy (grey curve), which we employ in this paper as the AGN quiescent state.
Figure 14: Radial profiles for the properties of an AGN thin disc, as predicted by the SS73 [61], SG03 [68], and TQM05 [69] models, as indicated (see the text for further details). Solutions for the gas density, aspect ratio, mid-disc temperature, and speed of sound are shown in the different rows of this plot array, from the upper to the lower panels, respectively. All solutions correspond to an accretion rate of $\dot{m}=0.025$ at the central engine, in Eddington units, and curves in different colours correspond to different assumed SMBH masses.
Figure 15: Spectral energy distributions predicted by the SS73, SG03, and TQM05 AGN disc models described in the text, as indicated. The grey curve represents the spectrum template for the host galaxy adopted in this paper as the quiescent AGN emission. The panels in different columns are obtained by assuming different accretion rates onto the central engine, whereas the panels in different rows correspond to different SMBH masses, as indicated.
Appendix B The initial pressure of the plasma ejection